OpenStudy (anonymous):

-5-(15y-1)=2(7y-16)-y I was absent and have no idea how to solve this

OpenStudy (anonymous):

What are we supposed to solve for/do?

OpenStudy (anonymous):

There is no equal sign so there really isn't a solution per se

OpenStudy (anonymous):

Solve for y, sorry I added a plus instead of an equal its fixed :)

OpenStudy (anonymous):

Oh ok. there we go.

OpenStudy (anonymous):

-5-15y+1 = 14y-32-y

OpenStudy (anonymous):

-15y - 4 = 13y - 32 -15y + 28 = 13y 28 = 28y 1 = y